### v1

# Exposing the Cracks: Vulnerabilities of Retrieval-Augmented LLM-based Machine Translation

Abstract
REtrieval-Augmented LLM-based Machine Translation (REAL-MT) shows promise for knowledge-intensive tasks like idiomatic translation, but its reliability under noisy retrieval contexts remains poorly understood despite this being a common challenge in real-world deployment. To address this gap, we propose a noise synthesis framework and new metrics to evaluate the robustness of REAL-MT systematically. Using this framework, we instantiate REAL-MT with Qwen-series models, including standard LLMs and large reasoning models (LRMs) with enhanced reasoning, and evaluate their performance on idiomatic translation across high-, medium-, and low-resource language pairs under synthesized noise. Our results show that low-resource language pairs, which rely more heavily on retrieved context, degrade more severely under noise than high-resource ones and often produce nonsensical translations. Although LRMs possess enhanced reasoning capabilities, they show no improvement in error correction and are even more susceptible to noise, tending to rationalize incorrect contexts. We find that this stems from an attention shift away from the source idiom to noisy content, while confidence increases despite declining accuracy, indicating poor calibration. To mitigate these issues, we investigate training-free and fine-tuning strategies, which improve robustness at the cost of performance in clean contexts, revealing a fundamental trade-off. Our findings highlight the limitations of current approaches, underscoring the need for self-verifying integration mechanisms.
Figure 1: Examples of correct and noisy contextual cues that may arise during online retrieval. The idiom and its translation are underlined. A robust REtrieval-Augmented LLM-based Machine Translation (REAL-MT) system should maintain fidelity in noisy scenarios.
Introduction
REtrieval-Augmented LLM-based Machine Translation (REAL-MT) is increasingly used to enhance translation quality for knowledge-intensive MT tasks like idiomatic translation (Li et al. 2024; Donthi et al. 2025). Although external knowledge can enhance translation performance, reliance on it is a double-edged sword: when the retrieved context contains noise such as irrelevant or misleading information, LLMs often produce nonsensical translations, as illustrated in Figure 1. Since noisy retrieval is unavoidable in real-world deployment, the behavior of REAL-MT under such conditions remains poorly understood, posing a critical barrier to deploying REAL-MT in safety-sensitive or real-world applications where reliability is paramount. This gap motivates our central question: to what extent does noisy retrieval compromise REAL-MT’s trustworthiness?
To investigate how noisy retrieval compromises REAL-MT’s trustworthiness, we develop a two-pronged approach: a controlled noise injection framework to simulate realistic retrieval failures, and specialized evaluation metrics to quantify their impact. While standard machine translation metrics like COMET (Rei et al. 2020) assess overall output quality, they fail to capture semantic fidelity in idiomatic translation and cannot distinguish whether errors stem from source misinterpretation or over-reliance on retrieved context. To address this gap, we propose two complementary metrics: Fidelity, which evaluates translation correctness with a focus on idiomatic accuracy, and Context Adoption Rate (CAR), which quantifies the extent to which models rely on external context. Together, they enable fine-grained analysis of both what the model gets wrong and why.
We evaluate REAL-MT systems instantiated with standard LLMs and large reasoning models (LRMs) across high-, medium-, and low-resource language pairs under controlled synthesized noise. Our results show that REAL-MT is far more vulnerable to knowledge-level errors, such as irrelevant or contradictory retrieval content, than to surface-level perturbations (e.g., word reordering). Moreover, the extent of performance degradation scales with the degree of semantic deviation: the more the retrieved context diverges from the intended meaning of idioms, the more severe the drop in translation fidelity. This sensitivity is especially pronounced in low-resource language pairs, as reflected in their higher CAR, which signal greater dependence on external context and leads to more severe degradation under noisy retrieval. Surprisingly, LRMs, despite their enhanced reasoning capabilities, show no improvement in error correction and are even more susceptible to noise, often rationalizing incorrect contexts. Across noise conditions, CAR remains consistently high even when retrieved content contradicts the source, prompting us to examine whether this reflects active model integration or coincidental similarity. Our attention analysis confirms the former: models consistently attend to retrieved context regardless of its correctness, demonstrating active integration. This uncritical reliance is further exacerbated by poor metacognitive awareness, as confidence increases despite declining accuracy, a sign of severe miscalibration and absent self-verification.
Given this uncritical reliance on retrieved context, we explore both training-free and fine-tuning strategies to improve REAL-MT robustness. While both enhance noise resistance, they incur a consistent trade-off: performance degrades under clean, accurate contexts. Fine-tuning yields better robustness overall but still fails to adjust reliance based on context quality. This persistent trade-off reveals that post-hoc mitigation cannot overcome the model’s fundamental inability to self-verify, underscoring the need for self-verifying integration mechanisms that validate retrieved content before adoption and enable rejection of noise while preserving accurate knowledge.
In summary, this work (1) introduces a controlled noise injection framework to evaluate REAL-MT robustness under realistic retrieval failures systematically, (2) proposes Fidelity and Context Adoption Rate (CAR) as fine-grained metrics to diagnose error sources in knowledge-intensive translation, and (3) evaluates training-free and fine-tuning mitigation strategies, revealing a fundamental trade-off between robustness under noise and performance in clean contexts due to uncritical reliance on retrieval and poor calibration, which underscores the need for self-verifying integration mechanisms in REAL-MT systems.
Related Work
Retrieval-Augmented LLM-based Machine Translation
Large language models (LLMs) have revolutionized machine translation (MT), especially for low-resource languages or domains where sufficient parallel corpora are lacking. However, when confronted with translation scenarios demanding specific background knowledge, relying solely on the LLM’s internal knowledge proves inadequate. In such instances, incorporating external knowledge to address the inherent limitations of LLMs becomes crucial (Merx et al. 2024; Chen et al. 2024; Zebaze, Sagot, and Bawden 2025). The prompt engineering capabilities of LLMs enable the integration of externally retrieved knowledge via prompting, without additional training. This makes prompt-based retrieval-augmented generation an efficient and flexible solution for MT. Recent studies by Li et al. (2024) and Donthi et al. (2024) leverage LLMs to enhance idiomatic translation by incorporating idiom-meaning pairs retrieved from offline knowledge bases directly into the prompt.
Prior work operates under the assumption that all introduced external knowledge is correct, an assumption that does not hold in real-world scenarios. This work systematically investigates the impact of introducing noisy context on translation systems. Given the increasing popularity of retrieval-augmented LLM-based MT, understanding LLM performance in the face of noisy input is crucial for enhancing translation system trustworthiness.
Robustness in Retrieval-Augmented Language Models
Retrieval-Augmented Language Models (RALM), reliant on external retrieved content, are susceptible to compromised reliability and robustness in their generated outputs when exposed to noise, irrelevant information, or malicious data  (Zhou et al. 2024; Park and Lee 2024; Shen et al. 2024; Yang et al. 2025). In response to the robustness challenges posed by RALMs, researchers have developed diverse approaches to enhance system robustness. Fang et al. (2024) introduces a named Retrieval-Augmented Adaptive Adversarial Training (RAAT) method to enhance the model’s ability to recognize and handle various types of noise.  Yoran et al. (2023) fine-tune the model using the parameter-efficient fine-tuning method QLoRA (Quantized Long-Range Attention) by using the synthetically generated noisy data to enhance its robustness to noisy data.  Xia et al. (2025) introduce a novel end-to-end self-reasoning framework that enhances the robustness, interpretability, and traceability of RALMs. This improvement is achieved by leveraging the reasoning trajectories generated by the LLMs themselves.
Departing from prior research predominantly focused on English-centric scenarios, this study presents the first systematic analysis of cross-lingual translation tasks that are explicitly designed to be non-English-centric.
Experimental Settings
Datasets
To analyze how resource availability affects the robustness of retrieval-augmented LLM-based machine translation (MT), we group translation directions into high-, medium-, and low-resource tiers following Joshi et al. (2020), based on parallel data availability and typological distance from English. We compile a dataset of idiomatic translations spanning ten language pairs, selected only from sources that provide reference translations or explicit semantic annotations, ensuring reliable interpretation of idioms and enabling controlled noise synthesis. The collection integrates multiple publicly available resources:
  • High-resource: IdiomsInCtx-MT (Stap et al. 2024) (English–German, German–English, Russian–English), the French, Japanese, and Korean idioms from Liu, Chaudhary, and Neubig (2023) and the KISS dataset [^1] https://github.com/Judy-Choi/KISS-Korean-english-Idioms-in-Sentences-dataSet, all paired with English;
  • Medium-resource: the Finnish–English idioms from Liu, Chaudhary, and Neubig (2023);
  • Low-resource: the Persian–English corpus from Rezaeimanesh, Hosseini, and Yaghoobzadeh (2025) (based on the PersianIdioms repository) and the Hindi–English corpus from Donthi et al. (2025).
This design enables fine-grained analysis of REAL-MT robustness across both resource levels and linguistic diversity.
Models
We select three representative models to cover different model sizes and reasoning modes: two standard LLMs, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct (Team 2024), and one large reasoning model (LRM), Qwen3-8B (Team 2025). Qwen3-8B uniquely supports seamless switching between thinking mode (enabled via enable_thinking=True) and non-thinking mode (enabled via enable_thinking=False), demonstrating significantly enhanced reasoning capabilities over prior Qwen instruct models. We include this LRM to investigate whether such advanced reasoning enables models to detect and reject noisy retrieval contexts during inference.
Following prior work that shows lower temperatures improve translation performance (Peng et al. 2023), we adopt greedy decoding (do_sample=False) to achieve both high output quality and reproducibility. Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct are evaluated with max_tokens=4096, while Qwen3-8B uses max_tokens=32768 following its official configuration to use the full context length. All experiments are run with a batch size of 40 using the vLLM (Kwon et al. 2023) framework on a single NVIDIA H800 GPU with 80GB VRAM.
Controlled Noise Context Generation
Type Text Sem. Var. Syn. Var.
IdiomIdiom kankkulan kaivoon - -
$\mathcal{G}$ down the drain Correct Correct
$\mathcal{N}_{\text{struct}}$ drain the down No Yes
$\mathcal{N}_{\text{literal}}$ to Kankkula’s well Incorrect (Relevant) No
$\mathcal{N}_{\text{semantic}}$ to Kankkula’s house Incorrect (Irrelevant) No
$\mathcal{N}_{\text{opposite}}$ to good use Incorrect (Opposite) No
Table 1: A case example illustrating the Finnish idiom: “kankkulan kaivoon”, its gold meaning ($\mathcal{G}$), and four types of generated noisy meanings. “Sem. Var.” denotes Semantic Variations, and “Syn. Var.” denotes Syntactic Variations.
Our analysis of online idiom dictionaries and search results shows that real-world retrieval failures often arise from incomplete phrase matching or morphological variations, which frequently lead to the retrieval of either unrelated information or literal interpretations of idioms. To faithfully simulate these knowledge-level errors, we design three levels of semantic noise that form a spectrum of increasing deviation from the correct idiomatic meaning:
Literal Translation ($\mathcal{N}_{\text{literal}}$): A word-by-word translation of the idiom.
Semantic-Perturbed Literal Meaning ($\mathcal{N}_{\text{semantic}}$): A variant of the literal meaning that maintains surface-level overlap but introduces a subtle semantic distortion.
Opposite Meaning ($\mathcal{N}_{\text{opposite}}$): An adversarial variant that directly contradicts the intended meaning of the idiom.
In addition, we include syntactic perturbations (e.g., word reordering) as a control condition to isolate the impact of knowledge-level errors from surface-level input variations:
Structure-Perturbed Gold Meaning ($\mathcal{N}_{\text{struct}}$): A syntactic variant of the gold meaning, with core semantics preserved. Together, these four noise types, summarized in Table 1, enable a systematic investigation of how noisy retrieval compromises REAL-MT’s trustworthiness.
To balance computational cost and coverage, we randomly sample 200 instances per translation direction and use the closed-source model gemini-flash-2.0 to generate noisy contexts based on carefully designed prompt templates (see Appendix A).
To validate that the synthesized noisy meanings align with the intended objectives, we conduct a quantitative analysis using the following lexical and semantic metrics:
Translation Edit Rate (TER): Measure the degree of structural perturbation by quantifying the edit operations required to align the gold meaning ($\mathcal{G}$).
Embedding Cosine Similarity (Sim): Measure the semantic similarity by computing the cosine similarity between the synthesized noisy meaning and both the gold meaning ($\mathcal{G}$) and the literal translation ($\mathcal{N}_{\text{literal}}$). We use the all-mpnet-base-v2  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 model, which is trained on above 1 billion sentences and shows strong performance on Semantic Textual Similarity (STS) tasks. Given its effectiveness in English-centric settings, it is suitable for our evaluation, where the target language is English.
Contradiction Rate (CR): Measure the percentage of the generated opposite meaning ($\mathcal{N}_{\text{opposite}}$) truly contradicts the gold meaning ($\mathcal{G}$). We use the NLI model: roberta-large-mnli (Liu et al. 2019) to classify the relationship between each pair. A higher rate indicates a more effective adversarial perturbation.
Across 10 language pairs, average TER($\mathcal{G}$, $\mathcal{N}_{\text{struct}}$) is 25.2, Sim($\mathcal{G}$, $\mathcal{N}_{\text{struct}}$) = 0.92, Sim($\mathcal{G}$, $\mathcal{N}_{\text{literal}}$) = 0.75, Sim($\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$) = 0.82, Sim($\mathcal{G}$, $\mathcal{N}_{\text{semantic}}$) = 0.73, and CR($\mathcal{G}$, $\mathcal{N}_{\text{opposite}}$) = 0.85, indicating that the generated noise aligns with the intended design. Full per-language results are presented in Table 1 in Appendix.
Pair Metric $r$ $\rho$ $\tau$
Fi→En COMET-22 0.4179 0.3891 0.3672
Fidelity 0.9194 0.9286 0.9091
Ja→En COMET-22 0.3191 0.3664 0.3266
Fidelity 0.8305 0.8196 0.7699
Fr→En COMET-22 0.5710 0.5428 0.5270
Fidelity 0.7822 0.7589 0.7378
Table 2: Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$ between human evaluations and automatic metrics.
Context HiEn FaEn FiEn JaEn FrEn KoEn RuEn DeEn EnFa EnDe $Avg_{C}$ $Avg_{F}$
[…]
Qwen2.5-7B-Instruct
$\mathcal{C}_{\text{none}}$ 0.8 65.8 0.6 53.2 0.4 59.5 1.1 60.2 1.5 77.6 1.6 73.8 1.8 71.9 1.7 61.3 0.8 64.5 1.5 65.3 1.2
$\mathcal{G}$ 2.1 78.4 2.5 66.0 2.2 67.2 2.4 67.3 2.5 79.0 2.6 78.9 2.7 80.1 2.7 63.3 1.1 75.2 2.2 72.8 2.3
$\mathcal{N}_{\text{struct}}$ 1.9 77.5 2.2 63.6 2.0 66.1 2.2 65.4 2.3 79.1 2.5 77.8 2.5 77.8 2.5 62.6 1.1 73.7 2.2 71.5 2.1
$\mathcal{N}_{\text{literal}}$ 1.3 64.9 1.1 55.4 0.7 62.7 1.5 58.6 1.5 77.6 1.9 73.7 1.9 69.5 1.6 59.3 0.9 72.4 2.0 66.0 1.4
$\mathcal{N}_{\text{semantic}}$ 0.8 63.3 0.8 53.0 0.5 60.8 1.2 58.4 1.4 76.5 1.4 73.2 1.6 67.8 1.3 57.2 1.4 69.1 1.4 64.4 1.2
$\mathcal{N}_{\text{opposite}}$ 0.3 70.9 0.5 57.0 0.4 61.4 0.9 59.7 0.7 77.0 1.4 73.3 1.2 66.7 0.8 63.3 0.7 68.6 1.2 66.4 0.8
Qwen2.5-14B-Instruct
$\mathcal{C}_{\text{none}}$ 1.3 72.6 1.2 57.4 0.6 66.6 1.7 65.1 2.1 79.8 2.0 77.5 2.1 75.22 1.9 67.2 1.2 77.2 2.0 71.0 1.6
$\mathcal{G}$ 2.3 82.2 2.6 67.5 2.4 68.3 2.4 67.2 2.6 80.1 2.7 80.2 2.7 80.8 2.8 70.7 1.5 76.8 2.4 72.8 2.3
$\mathcal{N}_{\text{struct}}$ 2.1 80.4 2.3 65.0 2.1 68.2 2.2 66.4 2.4 80.2 2.6 79.6 2.7 79.8 2.7 71.3 1.5 77.5 2.4 74.3 2.3
$\mathcal{N}_{\text{literal}}$ 1.5 69.0 1.2 56.7 0.7 64.5 1.6 61.2 1.7 79.1 2.1 55.6 2.1 72.5 1.8 65.7 1.1 75.1 2.1 66.6 1.6
$\mathcal{N}_{\text{semantic}}$ 0.9 67.3 1.0 54.7 0.5 61.8 1.3 57.3 1.2 78.0 1.7 73.0 1.6 68.3 1.2 61.1 0.8 69.5 1.5 65.7 1.2
$\mathcal{N}_{\text{opposite}}$ 0.5 74.8 0.5 58.7 0.5 62.2 0.9 59.3 0.6 78.3 1.3 74.1 1.1 67.5 0.7 67.1 0.7 71.0 0.9 68.1 0.8
Qwen/Qwen3-8B (non-thinking mode)
$\mathcal{C}_{\text{none}}$ 1.3 72.7 1.0 42.4 0.6 64.5 1.4 62.8 1.7 79.6 1.8 76.5 1.8 73.7 1.7 68.7 1.0 75.2 1.8 68.5 1.4
$\mathcal{G}$ 2.1 80.9 2.5 66.0 2.4 67.0 2.3 66.3 2.5 80.2 2.6 79.3 2.7 79.9 2.7 71.9 1.3 76.5 2.2 74.2 2.3
$\mathcal{N}_{\text{struct}}$ 2.0 79.7 2.2 62.4 2.0 65.9 2.1 64.4 2.3 80.0 2.5 78.2 2.6 78.1 2.4 72.1 1.5 76.3 2.2 73.0 2.2
$\mathcal{N}_{\text{literal}}$ 1.4 67.6 1.1 54.9 0.8 62.6 1.5 56.0 1.2 78.3 2.0 73.9 1.8 70.1 1.6 68.7 1.2 74.4 2.0 67.4 1.5
$\mathcal{N}_{\text{semantic}}$ 0.9 66.2 0.9 52.6 0.5 59.6 1.0 55.3 1.0 76.5 1.5 71.7 1.3 65.6 1.0 66.1 1.0 72.2 1.5 65.1 1.1
$\mathcal{N}_{\text{opposite}}$ 0.6 73.8 0.5 56.4 0.3 61.0 0.8 58.0 0.4 78.1 1.5 72.8 1.0 68.7 0.7 70.4 1.0 74.2 1.5 68.2 0.8
Qwen/Qwen3-8B (thinking mode)
$\mathcal{C}_{\text{none}}$ 1.3 70.3 1.0 55.1 0.6 64.6 1.4 62.6 1.8 79.8 1.9 75.9 1.8 74.5 1.7 70.7 1.2 77.4 2.0 70.1 1.5
$\mathcal{G}$ 2.2 81.9 2.5 66.5 2.4 66.2 2.5 66.1 1.2 80.3 2.7 78.9 2.7 79.7 2.8 74.3 1.5 79.8 2.3 74.9 2.3
$\mathcal{N}_{\text{struct}}$ 2.0 80.3 2.2 63.8 2.1 64.7 2.2 63.4 2.3 79.7 1.9 77.7 2.5 77.4 2.5 73.8 1.5 79.5 2.3 73.4 2.2
$\mathcal{N}_{\text{literal}}$ 1.4 67.2 1.1 55.4 0.8 61.3 1.4 54.7 1.5 77.5 0.7 72.9 1.8 68.2 1.5 67.4 1.1 76.2 2.0 66.8 1.3
$\mathcal{N}_{\text{semantic}}$ 0.6 65.2 0.8 53.3 0.5 57.2 0.8 54.0 0.9 74.5 1.2 70.2 1.2 63.9 1.0 62.7 0.9 68.7 1.2 63.3 0.9
$\mathcal{N}_{\text{opposite}}$ 0.1 73.6 0.5 56.9 0.2 57.9 0.3 56.0 0.1 76.3 0.6 70.8 0.5 65.0 0.2 68.9 0.5 70.0 0.6 66.2 0.4
Table 3: Performance of LLMs on idiom translation with various context settings, C denotes Comet-22, and F denotes Fidelity.
Metrics
Fidelity
Conventional machine translation metrics like BLEU (Papineni et al. 2002) and COMET (Rei et al. 2020) mainly measure lexical overlap or semantic alignment with reference translations. However, the key challenge in translating idioms is conveying their intended meaning. Therefore, we developed a metric specifically to assess meaning preservation in idiom translation, enabling more accurate evaluation of idiomatic translation quality.
Given a source sentence $x$ and retrieved context $c$, a model $M$ autoregressively generates a translation $y$. A “fidelity” score measuring how accurately the translation $y$ reflects the intended meaning $m$ of the source idiom. Following previous work (Liu et al. 2023; Li et al. 2024), we use a closed-source, high-performance language model gpt-4o-mini for the score’s evaluation. This score can be formatted as:
$\mathcal{F}(y,m)=\mathop{\arg\max}\limits_{r\in\{0,1,2,3\}}P(R=r\mid\text{Prompt}(y,m))$ (1)
where $\text{Prompt}(y,m)$ is a prompt function that generates a text prompt for the LLM, $R$ is a random variable representing the LLM’s output. To mitigate this uncertainty and improve the accuracy of the assessment, we integrate the LLM’s output over 20 runs. The fidelity score ranges from 0 to 3, where 0 indicates completely unfaithful and 3 indicates perfectly faithful. For detailed prompt information, please refer to Appendix B.1.
The REAL-MT system is robust if the value of $\mathcal{F}(y,m)$ does not decrease when the context $c$ is noisy. For instance, an LLM’s robustness is demonstrated if its translation remains unaffected even when the input meaning is incorrect.
Context Adoption Rate (CAR)
We design this metric to assess whether LLM utilizes the context $c$ we introduced for translation. When $c$ is noisy and the model adopts $c$, this indirectly indicates that the model lacks robustness. Given the variability in how different models translate idioms across languages and the resulting inconsistencies in fidelity quality, a single fidelity metric may not provide a clear and transparent measure of a model’s susceptibility to noise. Therefore, we introduce this metric to enhance the interpretability of how models process contextual information.
To calculate the CAR score, we first formalize the generation process of translation $y$ can be formalized as:
$P(y|x,c)=\prod_{i=1}^{n}P(y_{i}|y_{1},y_{2},...,y_{i-1},x,c)$ (2)
The generated translation $y={y_{1},y_{2},...,y_{n}}$, among them, each $y_{i}$ is the generated target word. The CAR score can be formally defined as:
$\text{CAR}(c,y)=\begin{cases}1,&\text{if }c\notin y_{\text{no\_context}}\land c\in y\\ 0,&\text{otherwise}\end{cases}$ (3)
where $y_{\text{no\_context}}$ is the translation without using context clues. We assign a score of 1 if the translation without context cues misses element $c$ while the translation using context cues successfully includes it; otherwise, a score of 0 is given. Therefore, when $c$ is noisy, the lower the CAR score, the more robust the model is. We leverage the high-performance language model gpt-4o-mini to perform this evaluation. The detailed prompt templates used for conducting evaluations are provided in Appendix B.2.
Human Evaluation
Prior work shows that reference-free metrics like CometKiwi (Rei et al. 2022b) struggle to capture idiomatic meaning (Li et al. 2024). To better evaluate idiom translation, we use the reference-based COMET-22 (Rei et al. 2022a) and introduce Fidelity as a complementary metric. We validate it via human evaluation, establishing a reliable ground truth and measuring its correlation with automatic scores. We evaluate on Fi→En, Ja→En, and Fr→En, representing diverse typological families and resource levels. Translations are generated using Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, and Qwen3-8B under the No Context ($\mathcal{C}_{\text{none}}$) setting, i.e., direct translation without external knowledge. For each language pair, the first 50 instances are annotated by three linguistics experts following the same guidelines as the models (see Appendix B.1). Scores are averaged to improve reliability. As shown in Table 2, Fidelity correlates highly with human judgments across languages, demonstrating that gpt-4o-mini can serve as an effective automatic evaluator. Given this strong alignment, we adopt Fidelity as the primary metric in subsequent experiments.
To assess CAR’s reliability, we conduct a human evaluation on 200 instances, achieving 72% accuracy. Given annotation costs, we use a cost-efficient LLM-based approach. Manual analysis shows errors are mostly false 0 rather than false 1. Therefore, the high CAR values we observe under noisy contexts provide a reliable lower bound on the LLM’s reliance, strengthening the validity of our findings.
To What Extent Does Noisy Retrieval Compromise REAL-MT’s Trustworthiness?
We evaluate REAL-MT across six retrieval conditions: No Context ($\mathcal{C}_{\text{none}}$), which measures performance without external knowledge, and Gold Meaning ($\mathcal{G}$), which provides an upper bound using oracle idiomatic knowledge, along with four noise variants, $\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$, $\mathcal{N}_{\text{opposite}}$, and $\mathcal{N}_{\text{struct}}$, that simulate realistic retrieval failures. Comparing performance across these settings enables a systematic analysis of how noise type and severity impact translation faithfulness.
Figure 2: Context Adoption Rate (CAR) of Qwen2.5-7B-Instruct in various contexts across ten language pairs.
Knowledge-level errors severely undermine REAL-MT’s trustworthiness unlike surface-level perturbations.
As shown in Table 3, models perform best under gold meaning ($\mathcal{G}$), confirming the benefit of correct context. Performance under syntactic perturbations ($\mathcal{N}_{\text{struct}}$) is comparable to $\mathcal{G}$, indicating that the model does not reject syntactically flawed text and possesses some syntactic self-correction capabilities.. With $\mathcal{N}_{\text{literal}}$, performance slightly improves in some low-resource languages (e.g., Hi→En, Fa→En), likely due to surface-level activation of idiom-related knowledge. However, when the context is irrelevant or conveys the opposite meaning ($\mathcal{N}_{\text{opposite}}$), translations increasingly align with the corrupted context rather than the source, resulting in a performance drop. These results show that semantic noise directly undermines the trustworthiness of REAL-MT, with worse performance as semantic distortion increases.
Large reasoning models rationalize rather than reason in thinking mode.
We observe that in thinking mode, models often produce reasoning traces contradicting the noisy context, indicating awareness of inconsistency, yet still generate outputs aligned with the noisy context (see Appendix Figure 4). This behavior suggests a tendency to rationalize rather than reason, potentially due to reward hacking during training, where models prioritize contextual coherence over factual fidelity (Chen et al. 2025). Supporting this, our quantitative analysis (Table 3) shows that Qwen3-8B in non-thinking mode consistently outperforms thinking mode under noisy contexts. For instance, under $\mathcal{N}_{\text{opposite}}$, non-thinking mode achieves an average Fidelity score of $0.8$, while thinking mode also drops to $0.4$. This disconnect between reasoning and output underscores the need for truth-preserving inference in large reasoning models.
Low-resource languages exhibit stronger reliance on retrieval context.
Figure 2 demonstrates that medium-to-low-resource language pairs (e.g., Hi→En, Fa→En, Fi→En) show significantly higher Context Adoption Rates (CAR) compared to high-resource pairs (e.g., En→De, En→Fr, De→En) across both clean ($\mathcal{G}$) and noisy conditions ($\mathcal{N}_{\text{struct}}$, $\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$, and $\mathcal{N}_{\text{opposite}}$). The pattern generalizes to other models as detailed in Appendix C. Such heightened reliance on contextual cues suggests that low-resource languages depend more heavily on external information during translation due to less parametric idiom knowledge. Consequently, LLMs become especially vulnerable to misleading information, even when the context conveys an opposite meaning, underscoring the critical need for robustness mechanisms in low-resource machine translation settings.
Figure 3: Attention allocation between source idiom and contextual meaning hint under various contexts.
Uncovering the Mechanism of Context Reliance in REAL-MT
Attention Shift from Idiom to Retrieved Context
LLMs exhibit high Context Adoption Rates (CAR) under noisy contexts, indicating that their outputs are strongly influenced by retrieved content, as shown in Figure 2. However, CAR only measures output similarity and does not reveal why the context shapes the generation process. To uncover the underlying mechanism, we analyze attention patterns during translation, as attention is widely used to diagnose where models focus during decoding (Wiegreffe and Pinter 2019). Specifically, we compute the average cumulative attention allocated to the source idiom versus the retrieved context across all target tokens.
As shown in Figure 3, attention consistently shifts toward the contextual meaning, even when it is incorrect or adversarial. This systematic pattern indicates active integration: the model consults the context during decoding, rather than merely producing aligned output. This confirms that the model’s predictions are anchored in the provided context, even when they semantically contradict the source input.
Language Mitigation $\mathcal{C}_{\text{none}}$ $\mathcal{G}$ $\mathcal{N}_{\text{struct}}$ $\mathcal{N}_{\text{literal}}$ $\mathcal{C}_{\text{semantic}}$ $\mathcal{N}_{\text{opposite}}$
Pairs Strategy Fidelity $\uparrow$ Fidelity $\uparrow$ CAR $\uparrow$ Fidelity $\uparrow$ CAR $\downarrow$ Fidelity $\uparrow$ CAR $\downarrow$ Fidelity $\uparrow$ CAR $\downarrow$ Fidelity $\uparrow$ CAR $\downarrow$
Fr→En Baseline 1.5 2.5 67% 2.3 69% 1.5 58% 1.4 61% 0.7 68%
Vanilla 1.9 2.4 44% 2.3 43% 1.5 56% 1.2 65% 1.1 49%
CDA 1.9 2.3 32% 2.3 31% 1.6 35% 1.6 35% 1.7 21%
ALI 1.9 2.2 30% 2.2 31% 1.8 28% 1.8 25% 1.8 11%
CK-PLUG 1.5 2.4 56% 2.2 53% 1.7 31% 1.6 39% 1.3 41%
Ja→En Baseline 1.1 2.4 75% 2.2 75% 1.5 60% 1.2 56% 0.9 64%
Vanilla 1.4 2.4 69% 2.3 67% 1.4 65% 1.1 70% 0.9 62%
CDA 1.4 2.0 51% 2.0 54% 1.6 48% 1.3 40% 1.4 23%
ALI 1.3 1.9 48% 1.8 47% 1.4 44% 1.3 36% 1.4 16%
CK-PLUG 1.1 2.1 65% 1.9 70% 1.6 57% 1.3 37% 1.3 32%
Hi→En Baseline 0.8 2.1 85% 1.9 79% 1.3 78% 0.8 80% 0.3 74%
Vanilla 0.8 2.2 81% 2.0 73% 1.3 72% 0.7 74% 0.4 65%
CDA 0.8 2.0 70% 1.7 61% 1.2 58% 0.8 62% 0.7 42%
ALI 0.8 1.7 54% 1.5 50% 1.1 52% 0.8 45% 0.8 26%
CK-PLUG 0.8 1.5 64% 1.6 57% 1.3 64% 1.1 71% 0.6 36%
Table 4: Performance of Qwen2.5-7B-Instruct after using different mitigation strategies, evaluated on three language pairs: Fr→En (high-resource), Ja→En (medium-resource), and Hi→En (low-resource).
Overconfidence in Context-Induced Errors
Given that REAL-MT consistently adopts retrieved context even when it is incorrect, we investigate the model’s confidence in these error-prone outputs. Specifically, under highly misleading noise ($\mathcal{N}_{\text{opposite}}$), does the model assign low confidence to its context-induced translations, signaling awareness of potential error? Such uncertainty would indicate good calibration and enable confidence-based self-correction or rejection mechanisms. Conversely, high confidence in erroneous outputs would reveal a dangerous overtrust in retrieval, undermining system trustworthiness. To assess this, we measure confidence using the entropy of the output probability distribution, with lower entropy indicating higher confidence. To ensure we evaluate the correct segment, we identify the target tokens corresponding to the source idiom through attention-based alignment: for each generated token, we compute its attention weights over the input and determine if it attends primarily to the idiom or contextual cue. The longest continuous sequence of idiom-aligned tokens is treated as the translated idiom span, and we report average entropy over this span.
As shown in Figure 4, the model exhibits increased confidence in the translation of idiom spans under the $\mathcal{N}_{\text{opposite}}$ setting, even though fidelity drops significantly. This inverse relationship reveals a critical failure in calibration: the model becomes more certain of its outputs despite their inaccuracy, suggesting that it implicitly treats the noisy context as authoritative and integrates it without sufficient verification.
Figure 4: Idiom-span translation confidence and Fidelity in Qwen2.5-7B-Instruct when adopting $\mathcal{N}_{\text{opposite}}$ context.
Mitigation Strategies
To counter REAL-MT’s blind trust in noisy retrieval and its lack of self-verification, we investigate training-free strategies for on-the-fly rejection of unreliable contexts and training-based methods to teach models to discern and downweight misleading knowledge.
Training-free Strategy
In RAG scenarios, prior work has explored the fusion of external knowledge to handle conflicts with internal knowledge. Given the similarity to REAL-MT, where noisy retrieved contexts may mislead translation, we investigate CK-PLUG (Bi et al. 2025), a training-free method that dynamically controls knowledge reliance based on context reliability. CK-PLUG computes Confidence Gain (CG) to measure the change in token-level entropy. For tokens with positive CG, which indicates the context is beneficial, the method blends the context-aware and internal distributions using $\alpha=0.5$. For tokens with negative CG, where the context may be harmful, external knowledge is fully suppressed. Although our analysis shows that REAL-MT exhibits poor calibration under noise, CG relies on relative confidence shifts rather than absolute certainty, potentially capturing whether the context stabilizes or disrupts the model’s internal prediction. We therefore evaluate CK-PLUG as a representative training-free strategy to test whether such entropy-based signals can still enable on-the-fly rejection of harmful contexts, despite overall miscalibration.
Training-based Strategy
To address REAL-MT’s tendency to blindly trust retrieved context even when it is misleading, we investigate training-based strategies that explicitly expose models to adversarial retrieval conditions during fine-tuning. Specifically, we construct training instances where the retrieved context conveys the opposite meaning of the source idiom, while the target translation remains correct. This encourages the model to learn to disregard misleading external knowledge and rely more on its internal representation when the two conflict. To prevent overcorrection (i.e., ignoring all context, including accurate ones), we also include noise-free conditions, No Context ($\mathcal{C}_{\text{none}}$) and Gold Meaning ($\mathcal{G}$) settings, during training. Our goal is to assess whether this fine-tuning scheme improves robustness to noisy contexts without sacrificing performance when retrieval is accurate.
Settings
We perform parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA) (Hu et al. 2022), with rank $r=16$ and scaling factor $\alpha=16$. Training is conducted for 50 epochs with a batch size of 2 and a learning rate of 2e-4. We use the AdamW optimizer with linear warmup and a cosine learning rate schedule. The experiments are conducted on a single NVIDIA H800 GPU with 80GB of VRAM.
Dataset
We select three language pairs, Fr→En, Ja→En, and Hi→En, with varying resource levels to evaluate the generalizability of the proposed strategy across different data scales. Following Liu, Chaudhary, and Neubig (2023), we use their released training sets, which contain 1,000, 1,456, and 507 sentence pairs, respectively. The test sets are the same as those used in the original study.
Models
We evaluate three fine-tuning strategies, each trained on a total dataset size equal to three times the original $\mathcal{C}_{\text{none}}$ set to ensure fair comparison:
[…]
This design directly addresses REAL-MT’s core failure mode: blind trust in retrieved context. By exposing models to misleading contexts with correct supervision (ALI), we encourage them to learn that not all retrieved knowledge is reliable. CDA further teaches discriminative reliance by promoting trust in accurate context ($\mathcal{G}$) and rejection of adversarial noise ($\mathcal{C}_{\text{none}}$), a crucial step toward self-verifying integration. Including clean conditions prevents overcorrection and preserves the performance when retrieval is accurate.
Results and Discussion
Confidence signals are unreliable for context filtering in REAL-MT.
As shown in Table 4, CK-PLUG yields only marginal robustness gains, revealing a fundamental flaw in entropy-based confidence signals: they assume that useful context reduces output entropy. Yet in low-resource REAL-MT, models are already overconfident in erroneous internal predictions. When accurate context ($\mathcal{G}$, e.g., Hi→En) contradicts this bias, it fails to lower entropy and may even increase uncertainty, causing CK-PLUG to suppress helpful knowledge. Rather than mitigating blind trust, this exacerbates the model’s reliance on flawed parametric knowledge, exposing the fragility of confidence-based methods in REAL-MT.
Training-based fine-tuning enables selective context reliance and outperforms training-free filtering.
[…]
Low-resource languages face greater difficulty in balancing noise robustness and performance with accurate retrieval.
Both training-free and training-based strategies improve robustness but degrade performance under $\mathcal{G}$, with the trade-off most pronounced in low-resource settings. This tension arises because neither approach can dynamically modulate reliance on retrieved context; instead, they apply fixed or static policies that inevitably sacrifice either noise resilience or clean-context utility. The strong dependence on hyperparameters like the blending weight $\alpha$ and training data proportions further underscores their brittleness. These findings reinforce the need for self-verifying integration mechanisms that can reliably assess context reliability at inference time, rejecting noise while preserving the benefits of accurate knowledge.
Conclusions
In this paper, we address the critical gap in understanding REAL-MT’s reliability under noisy retrieval, a common yet overlooked challenge in real-world deployment. To this end, we propose a controlled noise synthesis framework and two new metrics, Fidelity and Context Adoption Rate (CAR), to systematically evaluate REAL-MT’s robustness. Our evaluation, using Qwen-series models across high-, medium-, and low-resource language pairs, reveals that low-resource pairs are most affected by noise due to their stronger reliance on retrieval, often leading to nonsensical translations. Surprisingly, large reasoning models (LRMs) with enhanced reasoning capabilities not only fail to mitigate errors but are more vulnerable, frequently rationalizing incorrect contexts. This stems from a dual failure: attention shifts away from source idioms to noisy content, combined with rising confidence despite declining accuracy, highlighting poor calibration and weak self-verification. We explore both training-free and fine-tuning strategies to enhance robustness, but these come with trade-offs, such as reduced performance in clean contexts. Our findings highlight the limitations of current REAL-MT systems and emphasize the urgent need for self-verifying mechanisms to critically assess retrieved knowledge, allowing models to filter out noise while preserving the benefits of accurate external information.